OpenStudy (anonymous):

I have to evaluate the integral of x=0 to x=pi/3 of (1-sin(x))/cos(x)) du Te answer that i found is: ln|2+sqare root of 3| - ln|2| but the actual answer is ln(2+sqare root of 3). What did i do wrong?

OpenStudy (anonymous):

i guess you used u substitution for one of the expressions? what are you new limits of integration?

OpenStudy (anonymous):

Sec x - tanx

OpenStudy (anonymous):

\[\int\limits \frac{1-sinx}{cosx}dx= \int\limits (1-sinx)secxdx\]

OpenStudy (anonymous):

\[\int\limits (1-sinx)secxdx = \int\limits (secx-tanx)dx= \int\limits secx dx - \int\limits tanx dx\]

OpenStudy (anonymous):

\[\int\limits \tan(x) dx = - \ln(\cos(x))+c\]

OpenStudy (anonymous):

\[\int\limits \sec(x)dx = \ln(\tan(x)+\sec(x))+c\]

OpenStudy (anonymous):

\[ \ln(\cos(x))+\ln(\tan(x)+\sec(x))+c\]

OpenStudy (anonymous):

now u may plug in ur values to check the answers...